OpenStudy (anonymous):

Which equation is not equivalent to the formula d = rt? r equals t over d r equals d over t t equals d over r d = tr

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

d = rt d/t = r d/r = t these are all correct so what do you think the answer is ?

OpenStudy (anonymous):

The last one?

OpenStudy (texaschic101):

no....because d = tr is the same as d = rt the incorrect one is : r = t/d.....it is supposed to be r = d/t

OpenStudy (anonymous):

oh okay I got it, thank you very much.

OpenStudy (texaschic101):

no problem :)